### v1

@@ Abstract
Competitive programming problems increasingly serve as valuable benchmarks to evaluate the coding capabilities of large language models (LLMs) due to their complexity and ease of verification. Yet, current coding benchmarks face limitations such as lack of exceptionally challenging problems, insufficient test case coverage, reliance on online platform APIs that limit accessibility. To address these issues, we introduce LiveOIBench, a comprehensive benchmark featuring $403$ expert-curated Olympiad-level competitive programming problems, each with an average of $60$ expert-designed test cases. The problems are sourced directly from $72$ official Informatics Olympiads in different regions conducted between $2023$ and $2025$. LiveOIBench distinguishes itself through four key features: (1) meticulously curated high-quality tasks with detailed subtask rubrics and extensive private test cases; (2) direct integration of elite contestant performance data to enable informative comparison against top-performing humans; (3) planned continuous, contamination-free updates from newly released Olympiad problems; and (4) a self-contained evaluation system facilitating offline and easy-to-reproduce assessments. Benchmarking $32$ popular general-purpose and reasoning LLMs, we find that GPT-5 achieves a notable 81.76th percentile, a strong result that nonetheless falls short of top human contestant performance, who usually place above 90th. In contrast, among open-weight reasoning models, GPT-OSS-120B achieves only a 60th percentile, underscoring significant capability disparities from frontier closed models. Detailed analyses indicate that robust reasoning models prioritize precise problem analysis over excessive exploration, suggesting future models should emphasize structured analysis and minimize unnecessary exploration. All data, code, and leaderboard results will be made publicly available on our website.
@@ 1 Introduction
Coding has emerged as a critical domain for LLMs (Zhuo et al., 2024; Lai et al., 2022; Liu et al., 2024; Jimenez et al., 2024; Chan et al., 2024), with coding benchmarks serving as essential tools to evaluate LLMs’ algorithmic reasoning capabilities as these models continue advancing through inference-time scaling techniques (Li et al., 2022a; Kojima et al., 2023; DeepSeek-AI et al., 2025; OpenAI et al., 2024; Li et al., 2025b). However, rapid improvements in model capabilities have led to saturation of traditional coding benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), prompting the adoption of competitive coding benchmarks (Li et al., 2022a; Hendrycks et al., 2021b; Li et al., 2023; Shi et al., 2024) such as LiveCodeBench (Jain et al., 2024) and CodeELO (Quan et al., 2025), which leverage problems from platforms like Codeforces for their complexity and ease of verification. Despite their strengths, these benchmarks have notable weaknesses: (1) overestimation of LLMs’ performance due to high false-positive rates using incomplete test suites (Li et al., 2022a; Liu et al., 2023; Jain et al., 2024), (2) insufficient difficulty granularity and lacking exceptionally challenging questions (Jain et al., 2024; Quan et al., 2025), (3) usage of external APIs for evaluation, restricting reproducibility and accessibility (Jain et al., 2024; Quan et al., 2025; Zheng et al., 2025; Li et al., 2025c), (4) reliance on coarse pass rates as the sole evaluation metric, which misses the opportunity to gain insights on nuanced model capabilities (Jain et al., 2024; Li et al., 2022a; Wang et al., 2025; Shi et al., 2024), and (5) infrequent or costly updates due to the extensive human annotations and computational resources required (Wang et al., 2025; Zhu et al., 2025).
To address these gaps, we introduce LiveOIBench, the first comprehensive competitive coding benchmark constructed directly from Informatics Olympiads tasks, featuring expert-designed private tests, which will be made publicly available to support reproducible evaluation along with fine-grained scoring rubrics. Compared to previous benchmarks (Jain et al., 2024; Shi et al., 2024; Hendrycks et al., 2021b; Li et al., 2022a; Quan et al., 2025) and concurrent work (Li et al., 2025c; Zheng et al., 2025; Zhu et al., 2025; Wang et al., 2025) in Table 1, LiveOIBench features the following key advancements:
@@
    Expert-curated Tasks with Fine-grained Subtask Rubrics. We curate problems, test cases, and scoring rubrics directly from the official websites of $14$ Informatics Olympiads. This comprehensive test suite eliminates high false-positive rates common in previous benchmarks (Li et al., 2022a; Liu et al., 2023; Jain et al., 2024). Additionally, each task includes subtasks with scoring rubrics, enabling nuanced insights into model capabilities.
@@
In total, LiveOIBench comprises $403$ rigorously curated problems sourced from $72$ contests across $14$ Informatics Olympiads, each accompanied by an average of $60$ expert-written test cases. Using LiveOIBench, we evaluate $32$ leading models, revealing that proprietary models maintain a substantial performance advantage. In particular, GPT-5 (OpenAI, 2025b) achieves an average human percentile of $82$, while also exhibiting remarkable token efficiency by reaching this performance with fewer than 20K reasoning tokens, positioning it on the efficiency frontier (Figure 1). Among open-weight alternatives, Seed-OSS (ByteDance Seed Team, 2025) achieves the 54th percentile and Qwen3-32B (Yang et al., 2025b) reaches the 42nd percentile, both demonstrating significant performance gains from additional reasoning tokens. Additionally, GPT-OSS-120B (OpenAI et al., 2025) attains the 60th percentile, effectively narrowing the performance gap with GPT-5 and highlighting significant progress in open-weight model capabilities. Moreover, examining performance across different algorithms reveals current models’ weaknesses in algorithms like dynamic programming, which demand creative observation and hierarchical reasoning. Additionally, detailed reasoning trace analyses reveal that high-performing models strategically allocate more tokens to focused analysis rather than excessive exploration, underscoring that carefully managed reasoning behaviors are crucial for robust performance on challenging tasks.
@@
  • (Benchmarking Results) Conduct extensive benchmarking and detailed performance analysis of 32 leading open-source and proprietary models.
@@ 2 Related Work
The early code generation benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) mainly focus on the basic Python programs, which, for a long time, have been the standard ways to evaluate the code generation capability of LLMs. However, as the capability of LLMs evolves, simple benchmarks like HumanEval can no longer satisfy the benchmarking needs. Researchers have started developing more realistic and challenging benchmarks (Zhuo et al., 2024; Lai et al., 2022; Liu et al., 2024; Jimenez et al., 2024; Chan et al., 2024; Yin et al., 2023). Specifically, DS1000 (Lai et al., 2022) and ARCADE (Yin et al., 2023) consist of data science problems in Python. BigCodeBench (Zhuo et al., 2024) collects code generation tasks from Stack Overflow, which involves more complex instructions and diverse function calls. The SWE-Bench (Jimenez et al., 2024) takes one step further and tests models’ ability to solve real-world GitHub issues. This line of work emphasizes evaluating LLMs’ ability to effectively implement, debug, and reason through complex real-world coding tasks.
@@
In addition to real-world application benchmarks, there is another line of work: competitive programming benchmarks (Li et al., 2022a; Hendrycks et al., 2021b; Jain et al., 2024; Quan et al., 2025), which test the reasoning ability of models to solve challenging coding tasks within the specified time and memory constraints. All the previous competitive programming benchmarks collect problems from online coding platforms such as Codeforces and AtCoder, which do not release private test cases. The lack of sufficient private test cases may cause many false-positive solutions (Li et al., 2022a; Liu et al., 2023). Li et al. (2022a) augment test cases by mutating existing test inputs. Liu et al. (2023) leverages both LLM-based and mutation-based strategies to augment test cases with predefined rules. Even with over 200 additional tests per problem, Li et al. (2022a) shows there still exists nearly 50% false-positive rates. Other work (Quan et al., 2025; Zheng et al., 2025; Li et al., 2025c) tries to solve this problem by creating a platform to submit LLM-generated solutions directly to the Codeforces platform. Although this approach ensures that solutions are tested on the whole test set, its dependency on the online platform limits its accessibility to the research community, as large-scale evaluations involving thousands of submissions can overload platform servers.
@@
To solve the above problem, we collect problems from the official websites of many informatics Olympiads around the world. Most informatics Olympiads release their complete test set, which is curated carefully by the organizing committees. We are one of the first works to leverage problems from different informatics Olympiads and evaluate the models’ performance against human contestants. Prior research by Shi et al. (2024) exclusively used USACO problems with pass rate as the sole evaluation metric. Concurrent benchmarks, such as LiveCodeBench Pro (Zheng et al., 2025), HLCE (Li et al., 2025c), OI-Bench (Zhu et al., 2025), and AetherCode (Wang et al., 2025), also incorporate competitive programming tasks from sources like ICPC and IOI. However, LiveCodeBench Pro and HLCE primarily evaluate using Codeforces, limiting their accessibility. OI-Bench relies mostly on private, non-English school contests without continuous updates, while AetherCode uses LLM-generated tests and extensive human annotation with pass rate evaluation only. In contrast, our benchmark provides comprehensive coverage across diverse Olympiads, allows easy updates by directly collecting official test cases, and employs detailed evaluation metrics including subtask rubrics and human percentile comparisons.
@@ 3 LiveOIBench Construction
Competition Selection and Task Collection: We first curate a comprehensive list of globally recognized international Informatics Olympiads and selectively incorporate national contests from top-performing IOI countries where English task statements are available (See Table LABEL:tab:contest_dates). For each selected contest, we develop a custom crawler that systematically extracts English task statements (See Appendix A.5) directly from official competition websites, capturing details such as time and memory constraints, subtask specifications, test cases, official solutions, and contestant rankings. When official sites lack complete or up-to-date information, we supplement the data by retrieving missing details from established online platforms such as CSES [^1] https://cses.fi and LibreOJ [^2] https://loj.ac. To mitigate potential contamination from pre-training datasets, we strictly limit our dataset to contests held in 2023 and after. Additionally, we provide full descriptions of each competition along with official websites in Appendix A.6, ensuring selected contests have extensive historical data, consistent participant numbers, and regularly hosted events. Our benchmark will be continuously updated by leveraging monthly or annual problem releases from 14 actively maintained competition websites, allowing us to regularly expand our dataset with new contests and maintain an active leaderboard using a website similar to Figure A1.
Markdown Conversion and Quality Assurance: Given that many contests provide task statements exclusively as PDF documents, we employ Marker [^3] https://github.com/datalab-to/marker to automatically convert these PDFs into markdown format. We further utilize Gemini-2.0-Flash to automatically verify and correct these markdown texts. To ensure conversion accuracy, we manually inspect a sample of 40 tasks before batch processing. Additionally, we verify our evaluation judge and crawled test cases by executing the official solutions from contest organizers, using these solutions as the ground truth to confirm test-case correctness and the robustness of our evaluation judge.
@@
Ultimately, LiveOIBench comprises $403$ rigorously curated problems from $72$ competitions across $14$ Informatics Olympiads, conducted between $2023$ and $2025$. The benchmark statistics are detailed in Table A4, with a detailed description of our dataset construction methodology provided in Appendix A and competition information in Appendix A.6.
@@
  • Live Updates. Continuously updated with recent contests to minimize data contamination. All $14$ competitions described in Appendix A.6 in our benchmark will be updated.
@@ 4 Benchmarking Results
We evaluate a comprehensive set of $32$ LLMs. These models are categorized into three groups based on their accessibility and “thinking” capabilities: proprietary LLMs, open-weight thinking LLMs, and open-weight non-thinking LLMs. More details about models can be found Appendix B. During inference, we sample 8 candidate solutions per model and pick the solution with the highest score (Jain et al., 2024; Quan et al., 2025). We adopt the following evaluation metrics: pass rate (Kulal et al., 2019; Chen et al., 2021), relative score, human percentile, Olympics medal system, and Codeforces Elo (Quan et al., 2025; Zheng et al., 2025). With subtask rubrics and human contestants results, we can calculate each model’s total points in a contest, allowing precise comparisons to human contestants via percentile rankings and medal awards. The description of each metric can be found in Table 2 or Appendix C. In Table 2, we present benchmarking results for selected models that obtain the top performance in the corresponding categories of models. Full results for all evaluated models are included in Table A9.
@@
Table 2: Main results of best-performing models in each category evaluated on all $72$ contests. Full results for all 32 models are presented in Table A9. Gold and Medals: % of contests in which a model achieved a gold medal or any medal, respectively. Relative Score: % of total contest points obtained by the model. Human Percentile: % of human contestants that a model surpasses. Pass Rate: % of tasks where a model successfully passes all test cases. Elo: the Codeforces Elo rating earned by a model based on performance relative to human contestants. Higher is better for all metrics. Notably, the highest-performing GPT-5 achieves an impressive $81.76$th percentile but still falls short of top human contestants, successfully solving only 6363% of tasks in the benchmark.
Proprietary LLMs remain dominant, yet open-weight models are narrowing the performance gap. Our findings indicate that proprietary LLMs continue to lead in competitive coding benchmarks. Specifically, GPT-5 achieves impressive results, securing gold medals in $50\%$ of contests, winning medals of any type in $88.89\%$ of contests, and outperforming an average of $81.76\%$ of human contestants. Among open-source models tested, GPT-OSS-120B emerges as the strongest competitor. Under standard reasoning effort, GPT-OSS-120B achieves gold medals in $29.17\%$ of contests and performs near the 60th percentile—approximately $21.86$ percentile points below GPT-5. Notably, with high reasoning effort, GPT-OSS-120B surpasses Gemini-2.5-Pro and trails GPT-5 by merely $9$ percentile points. Seed-OSS, the second-best open-source model, attains the 54th percentile, narrowly trailing Gemini-2.5-Flash by only $3$ percentile points. However, other models exhibit substantial performance gaps, with Qwen3-32B and Deepseek-R1 obtaining gold medals in only $10\%$ and $7\%$ of contests, respectively, and performing at roughly the 4242nd percentile. Smaller and less powerful models, such as Qwen3-4B and DeepSeek-R1-Distill-Llama-8B, exhibit notably lower performance—Qwen3-4B secures gold medals in only $1.39\%$ of contests, while DeepSeek-R1-Distill-Llama-8B achieves no gold medals and ranks merely at the 33rd percentile. These results clearly demonstrate that achieving meaningful performance on competitive programming tasks in LiveOIBench requires LLMs with substantial reasoning capabilities.
@@
Apart from GPT-4.1 and Deepseek-V3, all other non-thinking models fail to exceed a $10\%$ pass rate, underscoring the critical importance of extended thinking in addressing complex competitive coding tasks. Extending this analysis, we investigate inference-time scaling techniques and find that both parallel  (Chen et al., 2021; Jain et al., 2024) and sequential  (DeepSeek-AI et al., 2025; Snell et al., 2024; Li et al., 2025a) scaling methods significantly enhance coding capabilities. In Figure 2, parallel scaling identifies maximum coding capacity but shows diminishing returns beyond a few attempts. While sequential scaling, by increasing the reasoning budget, allows smaller models to approach larger-model performance in Figure A4, reinforcing our earlier observation on the importance of extended thinking capabilities. For detailed analyses, see Appendix E.2.
@@
No evidence of temporal performance degradation. We examine the quarterly pass rates for four leading LLMs from Q1’$23$ through Q2’$25$ in Figure A3. Our analysis reveals no significant performance degradation coinciding with the models’ knowledge cutoffs, nor any signs of benchmark contamination. A more comprehensive analysis is provided in Appendix E.1.

### v2

@@ Abstract
Competitive programming problems increasingly serve as valuable benchmarks to evaluate the coding capabilities of large language models (LLMs) due to their complexity and ease of verification. Yet, current coding benchmarks face limitations such as lack of exceptionally challenging problems, insufficient test case coverage, reliance on online platform APIs that limit accessibility. To address these issues, we introduce LiveOIBench, a comprehensive benchmark featuring $403$ expert-curated Olympiad-level competitive programming problems, each with an average of $60$ expert-designed test cases. The problems are sourced directly from $72$ official contests of 14 Informatics Olympiads in different regions conducted between $2023$ and $2025$. LiveOIBench distinguishes itself through four key features: (1) meticulously curated high-quality tasks with detailed subtask rubrics and extensive private test cases; (2) direct integration of elite contestant performance data to enable informative comparison against top-performing humans; (3) planned continuous, contamination-free updates from newly released Olympiad problems; and (4) a self-contained evaluation system facilitating offline and easy-to-reproduce assessments. Benchmarking $34$ popular general-purpose and reasoning LLMs, we find that GPT-5 achieves a notable 81.76th percentile, a strong result that nonetheless falls short of top human contestants, who usually place above 90th. In contrast, among open-weight reasoning models, GPT-OSS-120B achieves only a 60th percentile, underscoring significant capability disparities from frontier closed models. Detailed analyses indicate that robust reasoning models prioritize precise problem analysis over excessive exploration, suggesting future models should emphasize structured analysis and minimize unnecessary exploration. All data, code, and leaderboard results are publicly available on our website.
@@ 1 Introduction
Coding has emerged as a critical domain for LLMs (zhuo2024bigcodebench; Lai2022DS1000; liu2023repobench; jimenez2024swebench; chan2024mle-bench), with coding benchmarks serving as essential tools to evaluate LLMs’ algorithmic reasoning capabilities as these models continue advancing through inference-time scaling techniques (alphacode; kojima2023largelanguagemodelszeroshot; deepseek-r1; openai2024openaio1card; li2025s). However, rapid improvements in model capabilities have led to saturation of traditional coding benchmarks such as HumanEval (chen2021codex) and MBPP (austin2021programsynthesislargelanguage), prompting the adoption of competitive coding benchmarks (alphacode; hendrycksapps2021; li2023tacotopicsalgorithmiccode; shi2024languagemodelssolveolympiad) such as LiveCodeBench (jain2024livecodebenchholisticcontaminationfree) and CodeELO (quan2025codeelobenchmarkingcompetitionlevelcode), which leverage problems from platforms like Codeforces for their complexity and ease of verification. Despite their strengths, these benchmarks have notable weaknesses: (1) overestimation of LLMs’ performance due to high false-positive rates using incomplete test suites (alphacode; evalplus; jain2024livecodebenchholisticcontaminationfree), (2) insufficient difficulty granularity and lacking exceptionally challenging questions (jain2024livecodebenchholisticcontaminationfree; quan2025codeelobenchmarkingcompetitionlevelcode), (3) usage of external APIs for evaluation, restricting reproducibility and accessibility (jain2024livecodebenchholisticcontaminationfree; quan2025codeelobenchmarkingcompetitionlevelcode; zheng2025livecodebenchproolympiadmedalists; li2025humanityscodeexamadvanced), (4) reliance on coarse pass rates as the sole evaluation metric, which misses the opportunity to gain insights on nuanced model capabilities (jain2024livecodebenchholisticcontaminationfree; alphacode; wang2025aethercodeevaluatingllmsability; shi2024languagemodelssolveolympiad), and (5) infrequent or costly updates due to the extensive human annotations and computational resources required (wang2025aethercodeevaluatingllmsability; zhu2025oibenchbenchmarkingstrongreasoning).
To address these gaps, we introduce LiveOIBench, the first comprehensive Informatics Olympiads coding benchmark constructed directly from official contest sources, featuring expert-designed private tests and official contestant rankings, which will be made publicly available to support reproducible evaluation along with fine-grained scoring rubrics. Compared to previous benchmarks (jain2024livecodebenchholisticcontaminationfree; shi2024languagemodelssolveolympiad; hendrycksapps2021; alphacode; quan2025codeelobenchmarkingcompetitionlevelcode) and concurrent work (li2025humanityscodeexamadvanced; zheng2025livecodebenchproolympiadmedalists; zhu2025oibenchbenchmarkingstrongreasoning; wang2025aethercodeevaluatingllmsability) in Table 1, LiveOIBench features the following key advancements:
@@
    Expert-curated Tasks with Fine-grained Subtask Rubrics. We curate problems, test cases, and scoring rubrics directly from the official websites of $14$ Informatics Olympiads. This comprehensive test suite eliminates high false-positive rates common in previous benchmarks (alphacode; evalplus; jain2024livecodebenchholisticcontaminationfree). Additionally, each task includes subtasks with scoring rubrics, enabling nuanced insights into model capabilities.
@@
In total, LiveOIBench comprises $403$ rigorously curated problems sourced from $72$ contests across $14$ Informatics Olympiads, each accompanied by an average of $60$ expert-written test cases. Using LiveOIBench, we evaluate $34$ leading models, revealing that proprietary models maintain a substantial performance advantage. In particular, GPT-5 (openai2025gpt5) achieves an average human percentile of $81.76$, while also exhibiting remarkable token efficiency by reaching this performance with fewer than 20K reasoning tokens, positioning it on the efficiency frontier (Figure 1). Among open-weight alternatives, Seed-OSS (seed2025seedoss36binstruct) achieves the 54th percentile and Qwen3-32B (yang2025qwen3technicalreport) reaches the 42nd percentile, both demonstrating significant performance gains from additional reasoning tokens. Additionally, GPT-OSS-120B (openai2025gptoss120bgptoss20bmodel) attains the 60th percentile, effectively narrowing the performance gap with GPT-5 and highlighting significant progress in open-weight model capabilities. Moreover, examining performance across different algorithms reveals current models’ weaknesses in algorithms like dynamic programming, which demand creative observation and hierarchical reasoning. Also, detailed reasoning trace analyses reveal that high-performing models strategically allocate more tokens to focused analysis rather than excessive exploration, underscoring that carefully managed reasoning behaviors are crucial for robust performance on challenging tasks. Finally, we conduct an in-depth analysis of data contamination in our benchmark, finding minimal correlation between time, task familiarity, and model performance.
@@
  • (Benchmarking Results) Conduct extensive benchmarking and detailed performance analysis of 34 leading open-source and proprietary models.
@@ 2 Related Work
The early code generation benchmarks such as HumanEval (chen2021codex) and MBPP (austin2021programsynthesislargelanguage) mainly focus on the basic Python programs, which, for a long time, have been the standard ways to evaluate the code generation capability of LLMs. However, as the capability of LLMs evolves, simple benchmarks like HumanEval can no longer satisfy the benchmarking needs. Researchers have started developing more realistic and challenging benchmarks (zhuo2024bigcodebench; Lai2022DS1000; liu2023repobench; jimenez2024swebench; chan2024mle-bench; yin-etal-2023-natural). Specifically, DS1000 (Lai2022DS1000) and ARCADE (yin-etal-2023-natural) consist of data science problems in Python. BigCodeBench (zhuo2024bigcodebench) collects code generation tasks from Stack Overflow, which involves more complex instructions and diverse function calls. The SWE-Bench (jimenez2024swebench) takes one step further and tests models’ ability to solve real-world GitHub issues. This line of work emphasizes evaluating LLMs’ ability to effectively implement, debug, and reason through complex real-world coding tasks.
@@
In addition to real-world application benchmarks, there is another line of work: competitive programming benchmarks (alphacode; hendrycksapps2021; jain2024livecodebenchholisticcontaminationfree; quan2025codeelobenchmarkingcompetitionlevelcode), which test the reasoning ability of models to solve challenging coding tasks within the specified time and memory constraints. All the previous competitive programming benchmarks collect problems from online coding platforms such as Codeforces and AtCoder, which do not release private test cases. The lack of sufficient private test cases may cause many false-positive solutions (alphacode; evalplus). alphacode augment test cases by mutating existing test inputs. evalplus leverages both LLM-based and mutation-based strategies to augment test cases with predefined rules. Even with over 200 additional tests per problem, alphacode shows there still exists nearly 50% false-positive rates. Other work (quan2025codeelobenchmarkingcompetitionlevelcode; zheng2025livecodebenchproolympiadmedalists; li2025humanityscodeexamadvanced) tries to solve this problem by creating a platform to submit LLM-generated solutions directly to the Codeforces platform. Although this approach ensures that solutions are tested on the whole test set, its dependency on the online platform limits its accessibility to the research community, as large-scale evaluations involving thousands of submissions can overload platform servers.
@@
To solve the above problem, we collect problems from the official websites of many informatics Olympiads around the world. Most informatics Olympiads release their complete test set, which is curated carefully by the organizing committees. We are one of the first works to leverage problems from different informatics Olympiads and evaluate the models’ performance against human contestants. Prior research by shi2024languagemodelssolveolympiad exclusively used USACO problems with pass rate as the sole evaluation metric. Concurrent benchmarks, such as LiveCodeBench Pro (zheng2025livecodebenchproolympiadmedalists), HLCE (li2025humanityscodeexamadvanced), OI-Bench (zhu2025oibenchbenchmarkingstrongreasoning), and AetherCode (wang2025aethercodeevaluatingllmsability), also incorporate competitive programming tasks from sources like ICPC and IOI. However, LiveCodeBench Pro and HLCE primarily evaluate using Codeforces, limiting their accessibility. OI-Bench relies mostly on private, non-English school contests without continuous updates, while AetherCode uses LLM-generated tests and extensive human annotation with pass rate evaluation only. In contrast, our benchmark provides comprehensive coverage across diverse Olympiads, allows easy updates by directly collecting official test cases, and employs detailed evaluation metrics including subtask rubrics and human percentile comparisons.
@@ 3 LiveOIBench Construction
Competition Selection and Task Collection: We first curate a comprehensive list of globally recognized international Informatics Olympiads and selectively incorporate national contests from top-performing IOI countries where English task statements are available (See Table LABEL:tab:contest_dates). For each selected contest, we develop a custom crawler that systematically extracts English task statements (See Appendix A.5) directly from official competition websites, capturing details such as time and memory constraints, subtask specifications, test cases, official solutions, and contestant rankings. When official sites lack complete or up-to-date information, we supplement the data by retrieving missing details from established online platforms such as CSES [^1] https://cses.fi and LibreOJ [^2] https://loj.ac. To mitigate potential contamination from pre-training datasets, we strictly limit our dataset to contests held in 2023 and after. Additionally, we provide full descriptions of each competition along with official websites in Appendix A.6, ensuring selected contests have extensive historical data, consistent participant numbers, and regularly hosted events. Our benchmark will be continuously updated by leveraging monthly or annual problem releases from 14 actively maintained competition websites, allowing us to regularly expand our dataset with new contests and maintain an active leaderboard.
Markdown Conversion and Quality Assurance: Given that many contests provide task statements exclusively as PDF documents, we employ Marker [^3] https://github.com/datalab-to/marker to automatically convert these PDFs into markdown format. We further utilize Gemini-2.0-Flash to automatically verify and correct these markdown texts. To ensure conversion accuracy, we manually inspect a sample of 40 tasks before batch processing. Additionally, we verify our evaluation judge and crawled test cases by executing the official solutions from contest organizers, using these solutions as the ground truth to confirm test-case correctness and the robustness of our evaluation judge. More details can be found in Appendix A.1.
@@
Ultimately, LiveOIBench comprises 403 rigorously curated problems from 72 contests across 14 Informatics Olympiads, conducted between $2023$ and $2025$. The benchmark statistics are detailed in Table A4, with a detailed description of our dataset construction methodology provided in Appendix A and competition information in Appendix A.6.
@@
  • Live Updates. Continuously updated with recent contests to minimize data contamination. All $14$ competitions described in Appendix A.6 will be updated periodically.
@@ 4 Benchmarking Results
We evaluate a comprehensive set of $34$ LLMs. These models are categorized into three groups based on their accessibility and “thinking” capabilities: proprietary LLMs, open-weight thinking LLMs, and open-weight non-thinking LLMs. More details about models can be found in Appendix B. During inference, we sample 8 candidate solutions per model in C++ and pick the solution with the highest score (jain2024livecodebenchholisticcontaminationfree; quan2025codeelobenchmarkingcompetitionlevelcode). We adopt the following evaluation metrics: pass rate (passk1; chen2021codex), relative score, human percentile, Olympiad medal system, and Codeforces ELO (quan2025codeelobenchmarkingcompetitionlevelcode; zheng2025livecodebenchproolympiadmedalists). With subtask rubrics and human contestants results, we can calculate each model’s total points in a contest, allowing precise comparisons to human contestants via percentile rankings and medal awards. The description of each metric can be found in Table 2 or Appendix C. In Table 2, we present benchmarking results for selected models that obtain the top performance in the corresponding categories of models. Full results for all evaluated models are included in Table A9.
@@
Table 2: Main results of best-performing models in each category evaluated on all $72$ contests. Full results are presented in Table A9. Gold and Medals: % of contests in which a model achieved a gold medal or any medal, respectively. Relative Score: % of total contest points obtained by the model. Human Percentile: % of human contestants that a model surpasses. Pass Rate: % of tasks where a model successfully passes all test cases. ELO: the Codeforces ELO rating earned by a model based on performance relative to human contestants. All metrics are higher the better. Notably, the highest-performing GPT-5 achieves an impressive $81.76$ percentile but still falls short of top human contestants, successfully solving only 6363% of tasks in the benchmark.
Proprietary LLMs remain dominant, yet open-weight models are narrowing the performance gap. Our findings indicate that proprietary LLMs continue to lead in competitive coding benchmarks. Specifically, GPT-5 achieves impressive results, securing gold medals in $50\%$ of contests, winning medals of any type in $88.89\%$ of contests, and outperforming an average of $81.76\%$ of human contestants. Among open-source models tested, GPT-OSS-120B emerges as the strongest competitor. Under standard reasoning effort, GPT-OSS-120B achieves gold medals in $29.17\%$ of contests and performs near the 60th percentile—approximately $21.86$ percentile points below GPT-5. Notably, with high reasoning effort, GPT-OSS-120B surpasses Gemini-2.5-Pro and trails GPT-5 by merely $9$ percentile points. Seed-OSS, the second-best open-source model, attains the 54th percentile, narrowly trailing Gemini-2.5-Flash by only $3$ percentile points. However, other models exhibit substantial performance gaps, with Qwen3-32B and Deepseek-R1 obtaining gold medals in only $10\%$ and $7\%$ of contests, respectively, and performing at roughly the 4242nd percentile. Smaller and less powerful models, such as Qwen3-4B and DeepSeek-R1-Distill-Llama-8B, exhibit notably lower performance—Qwen3-4B secures gold medals in only $1.39\%$ of contests, while DeepSeek-R1-Distill-Llama-8B achieves no gold medals and ranks merely at the 1111st percentile. These results clearly demonstrate that achieving meaningful performance on competitive programming tasks in LiveOIBench requires LLMs with substantial reasoning capabilities.
@@
Apart from GPT-4.1 and Deepseek-V3, all other non-thinking models fail to exceed a $10\%$ pass rate, underscoring the critical importance of extended thinking in addressing complex competitive coding tasks. Extending this analysis, we investigate inference-time scaling techniques and find that both parallel  (chen2021codex; jain2024livecodebenchholisticcontaminationfree) and sequential  (deepseek-r1; DBLP:journals/corr/abs-2408-03314; DBLP:journals/corr/abs-2502-14382) scaling methods significantly enhance coding capabilities. In Figure 2, parallel scaling identifies maximum coding capacity but shows diminishing returns beyond a few attempts. While sequential scaling, by increasing the reasoning budget, allows smaller models to approach larger-model performance in Figure A3, reinforcing our earlier observation on the importance of extended thinking capabilities. For detailed analyses, see Appendix E.2.